# Characterization of FBK NUV-HD-Cryo SiPMs near LHe temperature

### v3

@@ Abstract
Five FBK “NUV-HD-Cryo” SiPMs have been characterized at 7 K and 10 K, with 405 nm and 530 nm LED light, respectively. The dark current rate (DCR) was measured to be 1 Hz for the 100 mm2-size SiPMs, or 0.01 Hz/mm2, which is 7 orders lower than the DCR at room temperature (RT). Given the tiny DCR at these cryogenic temperatures, we measured the SiPMs’ I-V curves with such a method: illuminated the SiPMs with weak light, which differs from the conventional measurements at RT. Then, we measured the photo-detection efficiency (PDE), after-pulse (AP), and cross-talk (CT) with a bias voltage ranging from 6 to 11 V overvoltage (OV). At the OV interval (6 to 11 V), the PDE was between 20% - 45%, and the AP and CT were both between 5% and 20%. Suppose the bias is 10 V OV, the PDE would be 40%, and the AP and CT are 20%. Combining all of the measurements, we are confident that the SiPMs can be equipped as the photosensors on liquid helium detectors, including but not limited to the time projection chambers, which we have proposed in hunting for low-mass dark matter directly and beyond.
@@ I Introduction
Liquid helium (LHe) detectors have been proposed or employed to search for solar neutrinos [1, 2, 3], dark matter [4, 5, 6, 7, 8, 9, 10], and neutron electron dipole moment (nEDM) [11, 12, 13, 14]. When a particle interaction takes place within LHe, it will emit scintillation peaked at 80 nm [15, 16]. Given that no commercial photosensor is capable of detecting vacuum ultraviolet (VUV) light directly, a wavelength shifter such as tetraphenyl butadiene (TPB) is often implemented to convert it to visible light. As a rule of thumb, a photosensor should be as close as possible to the liquid or gas to avoid unnecessary light loss. Consequently, one of the requirements for the photosensor equipped on an LHe detector is that it should be functional near LHe temperature, 4.5 K 111For superfluid liquid helium detectors, the required temperature for a photosensor would be even lower than 4.5 K, and the exact temperature depends on the detector system.. Previous measurements have shown that photomultipliers (PMTs) and silicon photomultipliers (SiPMs) can work at 4.5 K or below [12, 18, 19, 20]. However, as far as we know, there is no complete characterization of a photosensor near 4.5 K. Although PMTs have been widely equipped on liquid argon and liquid xenon time projection chambers (TPCs) [21, 22, 23, 24, 25], they are not suitable for LHe TPCs because helium gas can permeate through the glass into the vacuum space of the photosensor. If that happens, an increase in dark current and degradation of the breakdown voltage is foreseen; as a result, the PMTs will deteriorate or even out of work since a discharge could happen inside the tube [26]. As such, SiPMs might be the only photosensor option for LHe TPCs.
In the manuscript, we present the characterization of Fondazione Brune Kessler (FBK) “NUV-HD-Cryo” SiPMs near LHe temperature, including an I-V curve, the dark count rate (DCR), photo-detection efficiency (PDE), and the after-pulse (AP) and cross-talk (CT) probabilities. The paper will be organized as follows. We first explain the testing setup in section II. The detailed tests of the I-V curve, DCR, PDE, AP, and CT at 10 K will be addressed in section III. After finishing the measurements at 10 K, we updated our detector, which made the SiPMs reach a simulation-consistent lower temperature of 7 K. We re-tested most of the parameters at 7 K again, as shown in section V. Our discussions and summaries are in section VI.
@@ II The testing setup introduction
ALETHEIA, standing for A Liquid hElium Time projection cHambEr In dArk matter, aims to hunt for low-mass dark matter (100’s MeV/c2 to 10 GeV/c2) with liquid helium-filled TPCs. To fully exploit the detector’s potential, the equipped photosensors should have a PDE greater than 40% near LHe temperature [9]. In addition, the I-V curve, DCR, AP, and CT must be measured to know whether the FBK NUV-HD-Cryo SiPMs’ performance suits the ALETHEIA project (and other applications mentioned above).
@@ II.1 The setup
As schematically shown in Fig. 1, the LED, the photodiode, and the SiPMs are all in a 10 K environment, while the preamplifier and data acquisition (DAQ) are at room temperature (RT). The 70 cm SMA cable going through the cryocooler’s walls connects the SiPMs in cryogenic conditions to the preamplifier on the warm side. All parts in cryogenic temperature have been verified earlier to ensure their capability of working at 4 K or below, and the parts were put in warm because they cannot work at LHe temperature.
@@
Generally, a preamplifier should be located as close as possible to a detector to reduce signal deterioration on the cable. Reference [27] used a preamplifier capable of working at 40 K, so the amplifier was (only) a few cm away from the SiPMs in cryogenic. High electron mobility transistors (HEMTs) have been implemented as a preamplifier at 4 K and 10 mK [28, 29]. In contrast, some other references [20, 30] put the preamplifier at RT. In our setup, since the TI OPA694 chip integrated into the preamplifier cannot be functional at LHe temperature, we must leave the preamplifier at RT. The preamplifier circuit is based on the ultra wide-band, low-power, current feedback operational amplifier, OPA694, and is configured as a non-inverting voltage amplifier with a gain of 360. The amplifier is connected to the ground through a low-pass RC filter, composed of a 20 serial resistor and a 1 nF capacitor, to reduce noise further. Depending on the data to take, an 8 GHz bandwidth, 25 GSa/s sample rate digital oscilloscope, or a Keithley 6485 electrometer will be utilized.
@@
Fig. 2 shows the actual setup; the cubic integrating sphere mounts on the barely visible cooling plate, which is the cooling source of the detector system. The LED, the photodiode, the SiPMs, and the temperature sensor (TS) are all installed on the surfaces of the sphere. Being driven by the pulses produced in a signal generator, DG535 222https://www.thinksrs.com/products/dg535.html, the LEDs can emit monochromatic light with an intrinsic FWHM of 5%. The “NUV-HD-Cryo” SiPMs, having an active window of 11.7 7.9 mm2, a cell pitch of 30 30 m2, was thankfully provided by FBK company. We bought the 10 10 mm2 (active area) photodiode from Opto diode company 333https://optodiode.com. For more info on the photodiode calibration, please refer to section II.2.
@@
Our team designed the integrating sphere and outsourced it to a commercial company in China. The inner surface of the 60 mm diameter integrating sphere is coated with highly reflective PTFE. The outside of the sphere is cubic, which gives convenience to (a) installing experimental parts on its surfaces and (b) being mounted securely on the cooling plate of the G-M cryocooler. As illustrated in Fig. 3, the LED, the calibrated photodiode, and the to-be-calibrated SiPMs are positioned 90 degrees from each other. This design ensures that the photodiode and the SiPMs receive the fixed ratio of photons as long as the size of the two corresponding ports does not change, regardless of the intensity profile of light emitted from the LED. In our case, the diameters of the ports for the photodiode and the SiPMs are 4.0 mm and 0.1 mm, respectively.
@@
The cooling plate of the G-M cryocooler approaches its lowest temperature after 3 hours of cooling; for the SiPMs on the integrating sphere, it will take a couple of additional hours to reach the lowest temperature. During our tests, we often run the cooling machine for more than 12 hours before launching a measurement to ensure the SiPMs and other parts on the sphere have reached the lowest temperature. Since it is impossible to measure the SiPMs’ temperature accurately, in the paper, the SiPMs’ temperature refers to the one measured on a TS 5 cm away. According to our simulation, the SiPMs’ actual temperature would be 0.3 K higher than the sensor-measured value, as will be further discussed in section IV. We did not test the SiPMs at a temperature higher than 10 K because it is not the typical temperature the photosensor equipped on LHe detectors is likely to work; neither were the SiPMs characterized below 7 K due to the cryocooler’s cooling power limitation.
@@ II.2 Photodiode responsivity calibration at NIM in China
As mentioned in references [33, 34, 35], a to-be-tested SiPMs’ PDE can be figured out by measuring the relative efficiency between the SiPMs and a calibrated photodiode. That being said, the efficiency (responsivity) of the photodiode should be characterized in advance. The photodiode under test is AXUV100G from Opto diode company. Although the company’s website has a datasheet, we cannot implement the responsivity data directly for a calibration near 4.5 K since the official data sheet was measured at RT. Our original plan was to calibrate a couple of diodes at the National Institute of Standards and Technology (NIST). However, NIST cannot do the calibration near LHe temperature for the moment, though they would be capable of doing that in the future [36]. Luckily, a team at the National Institute of Metrology (NIM) in China agreed to calibrate our photodiodes in their G-M cryocooler as a side-project of their primary research programs. Although the NIM facility can only calibrate a photodiode with 405 and 530 nm light, the calibration is still valuable since the two wavelengths line up well with TPB-shifted LHe scintillation, 350 - 600 nm [16].
@@ III.1 Testing the SiPMs’ I-V curve at 10 K
We tested the FBK SiPMs’ I-V curve at 4 K and RT in 2022 [9]. The I-V curves we measured at RT are consistent with the datasheet provided by FBK company at the same temperature. However, according to our communications with the company, they were unaware that any group has ever tested their SiPMs’ I-V curve near 4 K [37]. So, we were not sure whether our tests were correct or not in 2022, even though the shape of the I-V curve seemed different from the one at RT: when the bias voltage is less than a threshold, the current is nA; as long as the applied voltage reaches the criterion, the current jumps to mA quickly, as the blue curve in Fig. 5 shows. Although we repeated the I-V curve tests several times by different personnel, the same results were observed. The puzzle was solved in 2023, when the dark current rate (DCR) of the SiPMs was measured. Since the DCR was as low as 1 Hz at 10 K, the dark current (mainly composed of single electron events) was too small to build up a A current, even if the bias was higher than the breakdown voltage 444Assuming the SiPMs’ gain is 106 at 10 K with a bias higher than the breakdown voltage. Given the DCR is 1 Hz, the current would only be 0.16 pA = 1.6 10-19 C 1 Hz 106, where 1.6 10-19 C is the charge of an electron. At room temperature however, the DCR often is 107 Hz scale and the gain is 106, the current would be 1.6 A = 1.6 10-19 C 107 Hz 106.. As long as the bias voltage increases to a threshold much greater than 13 V overvoltage (OV), for example, the correlated noise increases significantly and eventually generates mA current [37]. The 1 Hz DCR is equivalent to 0.01 Hz/mm2 for the 100 mm2 SiPMs, which is 7 orders lower than the same photosensor at RT, 100 kHz/mm2 [27].
Given the especially low DCR near LHe temperature, we tested the SiPMs’ I-V curve in an alternative way: illuminating the SiPMs with weak pulsed light from an LED driven by a generator, DG535. During the DCR tests, the pulses used to drive the 405 nm LED were rectangular in shape, 4 V amplitude, 600 ns width, and 10 kHz frequency. Under such a testing setup, the I-V curve shows a conventional feature, as the green curve in Fig. 5. So far, we measured five randomly selected FBK SiPMs at 10 K with the same method, and all of the SiPMs recorded the same I-V curves. In addition, according to Fig. 5, the SiPMs’ breakdown voltage at 10 K is around 26.0 V, which is consistent with other tests [27, 24], where 27.0 V and 32.0 V breakdown voltage were obtained on the similar SiPMs (NUV-HD-LF) at 40 K and RT, respectively. It is observed that the lower the temperature, the lower the breakdown voltage, possibly because a lower temperature would result in a higher electron ionization rate, as predicted in reference [39].
@@ III.2 PDE measurements
The PDE of certain types of FBK SiPMs were measured at RT in references [34, 24]. A 40% PDE was obtained with the bias voltage of 5 V overvoltage (OV) or higher. However, according to our communications with FBK company, their clients have yet to fully characterize the SiPMs’ performance near LHe temperature [37]. Being inspired by the PDE calibrations on SiPMs performed at LAr or LXe temperatures [40, 41, 33, 42, 43, 34], we characterized the PDE of the NUV-HD-Cryo FBK SiPMs near LHe temperature with 405 nm and 530 nm light.

### v5

@@ Abstract
Five FBK “NUV-HD-Cryo” SiPMs have been characterized at 7 K and 10 K, with 405 nm and 530 nm LED light, respectively. The dark current rate (DCR) was measured to be 1 Hz for the 100 mm2-size SiPMs, or 0.01 Hz/mm2, which is 7 orders lower than the DCR at room temperature (RT). Given the very low DCR at these cryogenic temperatures, we measured the SiPMs’ I-V curves with such a method: illuminated the SiPMs with weak light, which differs from the conventional measurements at RT. Then, we measured the photo-detection efficiency (PDE), after-pulse (AP), and cross-talk (CT) with a bias voltage ranging from overvoltage (OV) 5 to 11 V. At the OV interval (5 to 11 V), the PDE was between 20% - 45%, and the AP and CT were both between 5% and 20%. With an OV higher than 10 V, the PDE would be 40%, and the AP and CT are 20%. Combining all of the measurements, we are confident that the SiPMs can be equipped as the photosensors on liquid helium detectors, including but not limited to the time projection chambers, which we have proposed in hunting for low-mass dark matter directly and beyond.
@@ I Introduction
Liquid helium (LHe) detectors have been proposed or employed to search for solar neutrinos [HERON1, HERON88, HERON96], dark matter [GuoMckinsey13, Maris17, SpiceHeRald21, Biekert22, ALETHEIA-EPJP-2023, QUEST-DMC23], and neutron electron dipole moment (nEDM) [Huffman2000, Ito2012, Ito2016, Phan20]. When a particle interaction takes place within LHe, it will emit scintillation light peaked at 80 nm [Kubota68, Benson18]. Given that no commercial photosensor is capable of detecting vacuum ultraviolet (VUV) light directly, a wavelength shifter such as tetraphenyl butadiene (TPB) is often implemented to convert it to visible light. As a rule of thumb, a photosensor should be as close as possible to the liquid or gas to avoid unnecessary light loss. Consequently, one of the requirements for the photosensor equipped on an LHe detector is that it should be functional near LHe temperature, 4.5 K 111For superfluid liquid helium detectors, the required temperature for a photosensor would be even lower than 4.5 K, and the exact temperature depends on the detector system.. Previous measurements have shown that photomultipliers (PMTs) and silicon photomultipliers (SiPMs) can work at 4.5 K or below [Ito2012, Cardini14, Iwai19, ZhangSiPM2022]. However, to the best of our knowledge, there is no complete characterization of a photosensor near 4.5 K. Although PMTs have been widely equipped on liquid argon and liquid xenon time projection chambers (TPCs) [LZ2022, XENONnT-NR-2023, PandaX2021, DarkSide20k17, DEAP2019], they are not suitable for LHe TPCs because helium gas can permeate through the glass into the vacuum space of the photosensor. If that happens, an increase in dark current and degradation of the breakdown voltage is foreseen; as a result, the PMTs will deteriorate or even out of work since a discharge could happen inside the tube [PMTsHamamatsu]. As such, SiPMs might be the only photosensor option for LHe TPCs.
In the manuscript, we present the characterization of Fondazione Brune Kessler (FBK) “NUV-HD-Cryo” SiPMs near LHe temperature, including a current-voltage (I-V) curve, the dark count rate (DCR), photo-detection efficiency (PDE), and the after-pulse (AP) and cross-talk (CT) probabilities. The paper will be organized as follows. We first explain the testing setup in section II. The detailed tests of the I-V curve, DCR, PDE, AP, and CT at 10 K will be addressed in section III. After finishing the measurements at 10 K, we updated our detector, which made the SiPMs reach a simulation-consistent lower temperature of 7 K. We re-tested most of the parameters at 7 K again, as shown in section V. Our discussions and summaries are in section VI.
@@ II The testing setup introduction
ALETHEIA, standing for A Liquid hElium Time projection cHambEr In dArk matter, aims to hunt for low-mass dark matter (100’s MeV/c2 to 10 GeV/c2) with liquid helium-filled TPCs. To fully exploit the detector’s potential, the equipped photosensors should have a PDE greater than 40% near LHe temperature [ALETHEIA-EPJP-2023]. In addition, the I-V curve, DCR, AP, and CT must be measured to know whether the FBK NUV-HD-Cryo SiPMs’ performance suits the ALETHEIA project (and other applications mentioned above). In previous works, the SiPMs made using the same technology have been fully tested at cryogenic temperatures, e.g. at liquid nitrogen temperature (77 K) and down to 40 K [Acerbi17, Capasso20, Razeto22, Acerbi2023]. For most of the parameters we tested in this work, we have seen comparable results (or in line with the decreasing trends with reducing temperature).
@@ II.1 The setup
As schematically shown in Fig. 1, the LED, the photodiode, and the SiPMs are all in a 10 K environment, while the preamplifier and data acquisition (DAQ) are at room temperature (RT). The 70 cm SMA cable going through the cryocooler’s walls connects the SiPMs in cryogenic conditions to the preamplifier on the warm side. All parts in cryogenic temperature have been verified earlier to ensure their capability of working at 4 K or below, and the parts were put in warm because they cannot work at LHe temperature. For the I-V curve tests, a Keithley 6485 electrometer was connected directly to the SiPMs; for the data-taking with the 8 GHz bandwidth, 25 GSa/s sample rate oscilloscope, an preamplifier was connected between the SiPMs and the oscilloscope.
@@
Generally, a preamplifier should be located as close as possible to a detector to reduce signal deterioration on the cable. Reference [Acerbi17] used a preamplifier capable of working at 40 K, so the amplifier was (only) a few cm away from the SiPMs in cryogenic. High electron mobility transistors (HEMTs) have been implemented as a preamplifier at 4 K and 10 mK [HEMTCryogenics19, HEMTCryogenics23]. In contrast, some other works [ZhangSiPM2022, Alessandro14] put the preamplifier at RT. In our setup, since the TI OPA694 chip integrated into the preamplifier cannot be functional at LHe temperature, we must leave the preamplifier at RT. We developed two versions of the preamplifier. The V1 amplifier is based on the ultra wide-band (690 MHz), low-power, current feedback operational amplifier, OPA694, and is configured as a non-inverting voltage amplifier. A low-pass RC filter, composed of a 20 serial resistor and a 1 nF capacitor is added to the operational amplifier output to reduce noise further. The V2 preamplifier has been updated to have a lower bandwidth of 170 kHz, a greater gain, and a better signal/noise ratio (SNR). The V1 was used in the measurements of DCR (Fig. 10) and AP (Fig. 11, Fig. 12, and Fig. 16). The V2 was implemented in the measurements of a raw signal (Fig. 6), a charge histogram (Fig. 7), PDE (Fig. 8 and Fig. 17), and CT (Fig. 13).
@@
Fig. 2 shows the actual setup; the cubic integrating sphere mounts on the barely visible cooling plate, which is the cooling source of the detector system. The LED, the photodiode, the SiPMs, and the temperature sensor (TS) are all installed on the surfaces of the sphere. Being driven by the pulses produced in a signal generator, DG535 222https://www.thinksrs.com/products/dg535.html, the LEDs can emit monochromatic light with an intrinsic FWHM of 5%. The NUV-HD-Cryo SiPMs, having an active area of 11.7 7.9 mm2, a cell pitch of 30 30 m2, was thankfully provided by FBK. We bought the 10 10 mm2 (active area) photodiode from Opto diode company 333https://optodiode.com. For more info on the photodiode calibration, please refer to section II.2.
@@
Our team designed the integrating sphere and outsourced it to a commercial company in China. The inner surface of the 60 mm diameter integrating sphere is coated with highly reflective PTFE. TPB has been implemented as a wavelength shifter below 2 K in other works [SpiceHeRald21, Biekert22, Ito2012, Ito2016, Phan20]. The outside of the sphere is cubic, which gives convenience to (a) installing experimental parts on its surfaces and (b) being mounted securely on the cooling plate of the G-M cryocooler. As illustrated in Fig. 3, the LED, the calibrated photodiode, and the to-be-calibrated SiPMs are positioned 90 degrees from each other (on different ports of the sphere). This design ensures that the photodiode and the SiPMs receive the fixed ratio of photons as long as the size of the two corresponding ports does not change, regardless of the intensity profile of light emitted from the LED. In our case, the diameters of the ports for the photodiode and the SiPMs are 4.0 mm and 0.1 mm, respectively.
@@
The cooling plate of the G-M cryocooler approaches its lowest temperature, 3 K, after 3 hours of cooling; for the SiPMs on the integrating sphere, it will take a couple of additional hours to reach the lowest temperature. During our tests, we often run the cooling machine for more than 12 hours before launching a measurement to ensure the SiPMs and other parts on the sphere have reached the lowest temperature. Since it is impossible to measure the SiPMs’ temperature accurately, in the paper, the SiPMs’ temperature refers to the one measured on a TS 5 cm away. According to our simulation, the SiPMs’ actual temperature would be 0.3 K higher than the sensor-measured value, as will be further discussed in section IV. We did not test the SiPMs at a temperature higher than 10 K because it is not the typical temperature the photosensor equipped on LHe detectors likely to work, though reference [Acerbi2023] tested the SiPMs between 75 K and 300 K; neither were the SiPMs characterized below 7 K due to the cryocooler’s cooling power limitation.
@@ II.2 Photodiode responsivity calibration at NIM in China
As mentioned in references [Eckert10, Zappala16, MPPCTechnicalNote21], a to-be-tested SiPMs’ PDE can be figured out by measuring the relative efficiency between the SiPMs and a calibrated photodiode. That being said, the efficiency (responsivity) of the photodiode should be characterized in advance. The photodiode under test is AXUV100G from Opto diode company. Although the company’s website has a datasheet, we cannot implement the responsivity data directly for a calibration near 4.5 K since the official data sheet was measured at RT. A typical option was to calibrate the photodiodes at the National Institute of Standards and Technology (NIST), but this did not happen because NIST do not offer such a calibration near LHe temperature for the moment, though they would be capable of doing that in the future [privateCommunicationsWithNIST]. Luckily, a team at the National Institute of Metrology (NIM) in China agreed to calibrate our photodiodes in their G-M cryocooler as a side-project of their primary research programs. Although the NIM facility can only calibrate a photodiode with 405 and 530 nm light, the calibration is still valuable since the two wavelengths line up well with TPB-shifted LHe scintillation, 350 - 600 nm [Benson18].
@@ III.1 Testing the SiPMs’ I-V curve at 10 K
We tested the FBK SiPMs’ I-V curve at 4 K and RT in 2022 [ALETHEIA-EPJP-2023]. The I-V curves we measured at RT are consistent with the measurements done in FBK laboratories at the same temperature. However, according to our communications with FBK, they were unaware that any group has ever tested their SiPMs’ I-V curve near 4 K. So, we were not sure whether our tests were correct or not in 2022, even though the shape of the I-V curve seemed different from the one at RT: when the bias voltage is less than a threshold, the current is nA; as long as the applied voltage reaches the threshold, the current jumps to mA quickly, as the blue curve in Fig. 5 shows. Although we repeated the I-V curve tests several times by different personnel, the same results were observed. The puzzle was solved in 2023, when the dark current rate (DCR) of the SiPMs was measured. Since the DCR was as low as 1 Hz at 10 K, the dark current (mainly composed of single electron events) was too small to build up a A current, even if the bias was higher than the breakdown voltage 444Assuming the SiPMs’ gain is 106 at 10 K with a bias higher than the breakdown voltage. Given the DCR is 1 Hz, the current would only be 0.16 pA = 1.6 10-19 C 1 Hz 106, where 1.6 10-19 C is the charge of an electron. At room temperature however, the DCR often is 107 Hz scale and the gain is 106, the current would be 1.6 A = 1.6 10-19 C 107 Hz 106.. As long as the bias voltage reaches a threshold much greater than the breakdown voltage (25 V), 38 V ( = 13 V OV), for example, the correlated noise increases significantly and eventually generates mA current. The 1 Hz DCR is equivalent to 0.01 Hz/mm2 for the 100 mm2 SiPMs, which is 7 orders lower than the same photosensor at RT, 100 kHz/mm2 [Acerbi17].
Given the especially low DCR near LHe temperature, we tested the SiPMs’ I-V curve in an alternative way: illuminating the SiPMs with weak pulsed light from an LED driven by a generator, DG535. During the I-V curve tests, the pulses used to drive the 405 nm LED were rectangular in shape, 4 V amplitude, 600 ns width, and 10 kHz frequency. Under such a testing setup, the I-V curve shows a conventional feature, as the green curve in Fig. 5. So far, we measured five randomly selected FBK SiPMs at 10 K with the same method, and all of the SiPMs recorded the same I-V curves. In addition, according to Fig. 5, the SiPMs’ breakdown voltage at 10 K is around 25.0 V, which is consistent with other tests [Acerbi17, DarkSide20k17], where 27.0 V and 32.0 V breakdown voltage were obtained on the similar SiPMs (NUV-HD-LF) at 40 K and RT, respectively. It is observed that the lower the temperature, the lower the breakdown voltage, possibly because a lower temperature would result in a higher electron ionization rate, as predicted in reference [Crowell68].
@@ III.2 PDE measurements
The PDE of certain types of FBK SiPMs were measured at RT in references [Zappala16, DarkSide20k17]. A 40% PDE was obtained with the bias voltage of 5 V overvoltage (OV) or higher. However, according to our communications with FBK, their clients have yet to fully characterize the SiPMs’ performance near LHe temperature. Being inspired by the PDE calibrations on SiPMs performed at LAr or LXe temperatures [Otte06, Finocchiaro09, Eckert10, Yang14, Piemonte16, Zappala16], we characterized the PDE of the NUV-HD-Cryo FBK SiPMs near LHe temperature with 405 nm and 530 nm light.